OpenStudy (anonymous):

log5(3x+1)+log5(x+1)=1 the (log5), 5 is a bit down the log :)

jhonyy9 (jhonyy9):

(3x+1)(x+1)/5

OpenStudy (anonymous):

start with \[\log_2((3x+1)(x+1))=1\] then \[(3x+1)(x+1)=5^1\]

jhonyy9 (jhonyy9):

1 is log5(5)

jhonyy9 (jhonyy9):

than coming on the left part will be with sign minus than result divide

OpenStudy (anonymous):

then \[3 x^2+4 x-4 = 0\] and then \[(x+2)(3x-2)=0\] so \[x=\frac{2}{3}\] and you can ignore the negative answer because you cannot take the log of a negative number

jhonyy9 (jhonyy9):

satellite why not is right ???

OpenStudy (anonymous):

@krypot \[\log_5(x)=y\iff x = 5^y\]

OpenStudy (anonymous):

*krypton, sorry

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so \[\log_5((3x+1)(x+1))=1\iff (3x+1)(x+1)=5\]

OpenStudy (anonymous):

am getting 2 as my answer

OpenStudy (anonymous):

yo am getting *2* as my answer

OpenStudy (anonymous):

got -2

jhonyy9 (jhonyy9):

3x2 +4x+1=0 x',2=(-4+/- sqrt(16-12))/6 =(-4+/- 2)/6 = -1 and - 1/3

jhonyy9 (jhonyy9):

how you have got this result ?

OpenStudy (anonymous):

i did log5(3x+1)/(x+1)=5 then cross multiplied it

OpenStudy (anonymous):

i got (3x+1)/(x+1)=5 i then cross mulitplied

jhonyy9 (jhonyy9):

no when we have loga + logb=log(a*b)

jhonyy9 (jhonyy9):

you need multiply not divide

OpenStudy (anonymous):

oh! thanks almost forgot my laws of indices.lol

jhonyy9 (jhonyy9):

bye and good luck